OpenStudy (anonymous):

Given the parent function f(x) = 2x, graph f(x) + 1.

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

there is no graphs

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@peachpi

OpenStudy (phi):

if you think of it as y= 2x and then add 1 y= 2x+1 that is the equation of a line. find two points, plot them , and draw a line through the two points.

OpenStudy (anonymous):

Alright. I think I understand. So, if I were to type this into desmos.com what would I do?

OpenStudy (ali2x2):

type in f(x) + 1 then press add slider i think

OpenStudy (anonymous):

ok let me try

OpenStudy (ali2x2):

https://www.desmos.com/calculator/cq94rv4dc8

OpenStudy (anonymous):

It didn't work.

OpenStudy (ali2x2):

there

OpenStudy (ali2x2):

press the play button

OpenStudy (anonymous):

@peachpi What r ur thoughts

OpenStudy (anonymous):

my thoughts are to type in the equation @phi gave

OpenStudy (anonymous):

kk. Thanks:) Can I ask another question?

OpenStudy (anonymous):

Given the parent function f(x) = 2^x, which graph shows f(x) + 1? Going back to the 1st question, the f(x)=2^x is a logarithmic ! So now how do I graph it??

OpenStudy (anonymous):

@phi

OpenStudy (phi):

y= 2^x is exponential when x is 0 you have y= 2^0 = 1 (any base to the zero power is 1) x=1, y= 2 x= 2, y= 4 x=3, y=8 plot those points, and put a smooth curve through them if you want y= 2^x + 1 , and 1 to each of the y values: (1,3) (2,5) (3,9) etc.  the curve looks roughly like